OpenStudy (anonymous):

What is the smallest possible value of f(x) if f(x)=Ѵ(3x-2)? A) 1 B) 0 C) 1/5

OpenStudy (anonymous):

f(x) ?

OpenStudy (anonymous):

if fx = root (3x -2)

OpenStudy (anonymous):

I don't get it.

OpenStudy (anonymous):

me neither

OpenStudy (anonymous):

The value under the radical must be positive. so 3x-2>=0 or x>=2/3. minimum is for x=2/3 so \( \sqrt{3x-2}=\sqrt{3\frac{2}{3}-2}=0 \)

OpenStudy (anonymous):